# Decoupling and Damping: Structurally-Regularized Gradient Matching for Multimodal Graph Condensation

### v1

@@ Appendix A APPENDIX
A.1. Analysis of Feature Alignment Effects on Multimodal Graph Condensation
We conduct multimodal graph condensation experiments with unaligned feature encoding to simulate real-world scenarios with feature inconsistency, where textual features are encoded using T5 (raffel2020exploring) (512-dimensional) and visual features using ViT (dosovitskiy2020image) (512-dimensional). This heterogeneous encoding creates semantic gaps in the feature space, providing a challenging environment to validate SR-GM’s robustness. As shown in Table 5, SR-GM maintains superior performance across all datasets and condensation ratios under unaligned conditions, achieving improvements of 7.81% over the sub-optimal method on Goodreads-NC at 0.05% condensation ratio and 3.11% on Amazon-Sports at 0.3% condensation ratio, confirming the effectiveness of gradient decoupling in mitigating multimodal gradient conflicts.
Comparative analysis between Table 2 (aligned features) and Table 5 (unaligned features) reveals that aligned conditions yield higher original dataset accuracy, with Ele-fashion achieving 85.68% versus 84.47%, Goodreads-NC reaching 79.45% versus 76.93%, and Amazon-Cloth attaining 71.59% versus 68.31%, reinforcing the crucial role of feature alignment in multimodal graph learning. However, the radar charts in Figure 5 demonstrate SR-GM’s distinctive advantage: in both aligned and unaligned scenarios, it exhibits the smallest performance gap relative to original datasets, with Figure 5 particularly showing significantly smaller performance degradation compared to other condensation algorithms under unaligned conditions.
A.2. Ablation Study on Gradient Decoupling and Structure Regularization
@@
The ablation study results fully validate the effectiveness of the SR-GM framework in addressing gradient conflict issues in multimodal graph condensation. In feature-aligned CLIP encoding scenarios, structure regularization (R) demonstrates significant advantages, particularly on the Goodreads-NC dataset where the $\Delta_{2}$ increments reach +7.56 to +8.88. This confirms our theoretical hypothesis regarding “structural damping” - by constraining the gradient field through Laplacian regularization, it effectively suppresses the propagation and amplification of local gradient conflicts within the graph structure. However, in high condensation ratio scenarios on the Ele-fashion dataset, the method relying solely on structure regularization ($\Delta_{1}=-1.18$) shows limitations, reflecting that in feature-aligned and data-sufficient environments, structural constraints alone may be insufficient to completely resolve the deep optimization problems caused by inter-modal semantic gaps.
@@
In contrast, the core value of gradient decoupling (D) becomes prominent in feature-unaligned T5VIT encoding environments. The stable $\Delta_{1}$ increments of +2.58 to +4.00 on the Ele-fashion dataset directly confirm the crucial role of the gradient decoupling mechanism in handling conflicts among multimodal gradient signals. Particularly noteworthy is the negative $\Delta_{2}$ increment observed for structure regularization under T5VIT encoding, a phenomenon that profoundly reveals how traditional structural constraints might instead reinforce incorrect gradient propagation paths in feature-unaligned environments. Gradient decoupling, through orthogonal projection to eliminate inter-modal gradient contradictions, provides a cleaner optimization foundation for structure regularization.
@@
These findings strongly support the dual-design philosophy of SR-GM: gradient decoupling resolves multimodal gradient conflicts at their source, while structure regularization suppresses conflict amplification along propagation paths. The synergistic effect of these two components enables SR-GM to maintain robust performance across diverse multimodal graph scenarios, providing reliable technical support for practical applications of multimodal graph condensation.

### v2

@@ 3.3. Analysis of Feature Alignment Effects on Multimodal Graph Condensation (Q2)
3.4. Condensation Efficiency (Q3)
We evaluate the computational efficiency of SR-GM by comparing its convergence behavior and training latency against the GCond. Specifically, we analyze the convergence trajectory of test accuracy relative to training epochs and assess the potential overhead introduced by our method.
As illustrated in Fig. 5, SR-GM converges faster by 220 and 50 epochs on the Ele-fashion and Goodreads-NC datasets, respectively. And Table 4 indicates that the average per-epoch training time of SR-GM remains nearly identical to that of GCond. These results demonstrate that SR-GM not only achieves superior performance but also maintains a competitive condensation speed.
3.5. Ablation Study on Gradient Decoupling and Structural Regularization (Q4)
@@
The ablation study results in Table 5 fully validate the effectiveness of the SR-GM framework in addressing gradient conflict issues in multimodal graph condensation. In feature-aligned CLIP encoding scenarios, structural regularization (R) demonstrates significant advantages, particularly on the Goodreads-NC dataset where the $\Delta_{2}$ increments reach +7.56 to +8.88. This confirms our theoretical hypothesis regarding “structural damping” - by constraining the gradient field through Laplacian regularization, it effectively suppresses the propagation and amplification of local gradient conflicts within the graph structure.
@@
However, in feature-unaligned T5-ViT encoding environments on the Ele-fashion dataset, the method relying solely on structural regularization ($\Delta_{2}=-3.44$) shows limitations, reflecting that structural constraints alone may be insufficient to completely resolve the deep optimization problems caused by inter-modal semantic gaps. In contrast, the core value of gradient decoupling (D) becomes prominent. The stable $\Delta_{1}$ increments of +2.58 to +4.00 on the Ele-fashion dataset directly confirm the crucial role of the gradient decoupling mechanism in handling conflicts among multimodal gradient signals. Particularly noteworthy is the negative $\Delta_{2}$ increment observed for structural regularization under T5-ViT encoding, a phenomenon that profoundly reveals how traditional structural constraints might instead reinforce incorrect gradient propagation paths in feature-unaligned environments. Gradient decoupling, through orthogonal projection to eliminate inter-modal gradient contradictions, provides a cleaner optimization foundation for structural regularization.
@@
These findings strongly support the dual-design philosophy of SR-GM: gradient decoupling resolves multimodal gradient conflicts at their source, while structural regularization suppresses conflict amplification along propagation paths. The synergistic effect of these two components enables SR-GM to maintain robust performance across diverse multimodal graph scenarios, providing reliable technical support for practical applications of multimodal graph condensation.
3.6. Ablation Study on the Impact of Multimodal Features (Q5)
To investigate the impact of multimodal features on the performance of graph condensation, we test the performance of GCond when condensing three types of input features: text features, image features and fused text-image features. Then we compare it against SR-GM when condensing fused features.
The node classification accuracy at Ele-fashion and Goodreads-NC is presented in Fig. 6. The fused multimodal data reduces the performance of GCond by approximately 6%. This may due to the inherent conflict between different modalities. In contrast, SR-GM outperform GCond by 14% at Ele-fashion and 9% at Goodreads-NC. It shows SR-GM’s ability to handle the multimodal graph data.
Figure 6. Test accuracy (%) for SR-GM and GCond using different input features: text-only, image-only, and fused text-image features.
4. Conclusion
Appendix A APPENDIX
A.1. Related Work
A.1.1. Graph Dataset Condensation.
Dataset Condensation (DC), also known as Dataset Distillation, aims to synthesize a small, informative dataset from a large original one such that models trained on the synthetic set achieve performance comparable to those trained on the original data. This paradigm was first pioneered in the image domain, where methods were developed to optimize synthetic pixels directly through a bi-level optimization framework (Wang et al., 2018). The core idea is to match certain properties between the real and synthetic data. Various matching criteria have been proposed, including matching model performance (Zhao et al., 2020) and feature distributions (Zhao and Bilen, 2023; Wang et al., 2022). Among these, matching the gradients of model parameters has been proven to be a particularly effective strategy (Zhao et al., 2020). The extension of these techniques to the graph domain, termed Graph Condensation (GC), is a nascent and challenging endeavor. Jin et al. (Jin et al., 2021) first extended DC to graphs, introducing a node-level condensation method. Unlike image data, graph data involves an intricate coupling between node attributes and topological structure, making the joint synthesis of features and edges a complex optimization task. Current GC methods primarily focus on preserving information in single-modal graphs, exploring strategies such as gradient matching (Jin et al., 2021), training trajectory matching (Zheng et al., 2023), and distribution matching (Liu et al., 2022). Recent research has evolved from developing efficient and interpretable graph condensation methods (Fang et al., 2024; Wang et al., 2024) to creating specialized techniques tailored for specific graph types, such as heterogeneous graphs and signed graphs (Gao et al., 2024; Li et al., 2025). However, these methods largely overlook the unique challenges posed by multimodal graphs, where node features themselves are composed of heterogeneous information from sources like text and images. The problem of inter-modal signal conflict within the GC framework remains a critical and underexplored gap. Our work aims to bridge this gap.
A.1.2. Multimodal Graph Learning.
Multimodal Graph Learning (MGL) has emerged as a vital field for analyzing complex real-world systems where entities are described by multiple data types (e.g., text, images, audio). MGL aims to leverage the relational structure of graphs to fully explore associations both within and between different modalities. This is particularly challenging because data fusion must be guided by the graph topology, which itself may vary across modalities. A key challenge in MGL is bridging the ”semantic gap”, where different modalities may provide conflicting or inconsistent representations of the same entity (Zhu et al., 2025; Liang et al., 2022). For instance, two products might be visually similar but have different textual descriptions, creating discrepancies that unimodal models cannot resolve. To address this, significant research has focused on feature alignment, which aims to map representations from different modalities into a unified embedding space. Contrastive learning frameworks, such as CLIP (Radford et al., 2021), have been notably successful in learning aligned vision-language representations. From an architectural perspective, methods like Multimodal Graph Convolutional Networks (MGCN) (Wei et al., 2019) and Multimodal Graph Attention Networks (MGAT) (Tao et al., 2020) have been proposed to explicitly model and fuse information from different modalities. While these MGL methods focus on designing more sophisticated Graph Neural Network (GNN) architectures to better handle existing multimodal data, they do not address the data-centric challenge of scalability (Ning et al., 2025; Cai et al., 2024). The question of how to effectively compress large-scale multimodal graphs, particularly in a way that handles the feature conflicts inherent in the condensation process, remains largely unaddressed. Our work bridges this gap by introducing a condensation method specifically designed with an optimization scheme attuned to the uniqueness of multimodal graphs.
A.2. Complexity Analysis
We analyze the computational complexity of the proposed SR-GM framework, focusing on the additional overhead introduced by the gradient decoupling and structural regularization modules.
A.2.1. Gradient Decoupling Complexity
The gradient decoupling process operates primarily on the gradient vectors of the synthetic nodes. As shown in the Algorithm 1, it involves computing element-wise dot products, norms, and vector projections for the text and image modalities. These operations are linear with respect to the number of synthetic nodes and do not involve node-to-node interactions. For $N^{\prime}$ synthetic nodes with $d$-dimensional features, the complexity of this component is $\mathcal{O}(N^{\prime}d)$.
A.2.2. Structural Regularization Complexity
The structural regularization minimizes the Dirichlet energy of the decoupled gradient field, formulated as Eq. (13). The computational bottleneck lies in the matrix multiplication, which has a complexity of $\mathcal{O}((N^{\prime})^{2}d)$. The subsequent trace operation takes $\mathcal{O}(N^{\prime}d)$. Thus, the complexity is dominated by $\mathcal{O}((N^{\prime})^{2}d)$.
A.2.3. Overall Time Complexity
Combining the two components, the total additional complexity is dominated by $\mathcal{O}((N^{\prime})^{2}d)$. It is important to note that the size of the condensed graph $N^{\prime}$ is extremely small. Consequently, the term $(N^{\prime})^{2}$ remains negligible compared to the complexity of training GNNs on the original large-scale graph. Therefore, SR-GM improves condensation quality with marginal computational overhead.
A.3. Generalizability Analysis (Q6)
To demonstrate that the effectiveness of our proposed SR module is not confined to a condensation paradigm or neural architecture, we investigate the generalizability of SR-GM from three perspectives: (i) compatibility with different matching objectives, (ii) transferability of the condensed graphs across various GNN backbones, and (iii) the performance stability when employing different architectures during the condensation process.
Table 6. Performance and efficiency comparison between GCDM and SR-GCDM. All datasets use CLIP encoding. Performance is reported as test accuracy (%) for Ele-fashion, Goodreads-NC and AUC score for Amazon-Sports, Amazon-Cloth. Total time required for condensation to converge is measured in seconds. The better accuracy is bolded.
Accuracy Time(s)
Dataset Ratio ($r$) GCDM SR-GCDM GCDM SR-GCDM
[…]
A.3.1. Generalization to Distribution Matching
To evaluate the versatility of the SR module, we extend its application to Distribution Matching (DM). Specifically, we integrate the SR module into GCDM, denoted as SR-GCDM.
Table 6 summarizes the performance and efficiency comparison between the original GCDM and our SR-enhanced version. Specifically, SR improves GCDM’s performance on most datasets, with gains of approximately $3\%$ on Ele-fashion and $10\%$ on Goodreads-NC. Similar to our observations with SR-GM, while the incorporation of the SR module introduces a slight computational overhead, the increase in total time required for condensation to converge is marginal compared to the performance boost. This demonstrates that SR is a plug-and-play component that is agnostic to the specific matching objective.
Dataset Methods GCN MLP MMGCN GAT SAGE Avg.
Ele-fashion GCond 78.87 74.94 71.53 34.65 64.46 64.89
$r=0.1\%$ SR-GM 79.91 81.71 75.88 39.31 81.50 71.66
Goodreads-NC GCond 62.07 49.06 43.81 33.41 53.56 48.38
$r=0.05\%$ SR-GM 67.76 57.34 51.36 34.33 64.16 54.99
A.3.2. Cross-Architecture Transferability
Condensed graphs should exhibit transferability across diverse GNN architectures, regardless of the model used during the condensation. We condense the datasets using SR-GM with a GCN in the condensation process and subsequently evaluate the performance of the condensed graphs on several unseen architectures.
As shown in Table 7, we test GraphSAGE (Hamilton et al., 2017), GAT (Velickovic et al., 2017), a multimodal GNN (MMGCN) (Wei et al., 2019), and a simple MLP (Taud and Mas, 2017). And we choose Ele-fashion and Goodreads-NC to report the performance. The graphs condensed by SR-GM show strong transfer performance on other architectures, maintaining high accuracy when the downstream model differs from the condensation model.
Table 8. Test accuracy (%) of SR-GM on the Ele-fashion datasets when using different backbone models for both condensation and evaluation. ”SAGE” denotes GraphSAGE. ”Original” means training on original graph. The best results are highlighted in bold.
Backbone Ratio ($r$) SR-GM GCond Original
[…]
A.3.3. Architectural Agnosticism of the Condensation Process
We investigate whether the condensation performance of SR-GM is sensitive to the choice of the GNN model used during the matching process. As shown in Table 8, we report the performance at Ele-fashion. For each run, the same model (e.g., GCN) is used for condensation and downstream evaluation.
From the table, we observe that SR-GM consistently constructs effective condensed graphs across all tested architectures. This validates that SR-GM does not need a specific GNN architecture to achieve superior condensation results.
A.4. Analysis of Regularization Weight $\lambda$ (Q6)
A critical hyperparameter in SR-GM is the regularization weight, denoted by $\lambda$. To investigate its impact on the graph condensation process, we condense graphs with various settings of $\lambda$ and condensation ratio, then report the accuracies in Fig. 7. The darker shades indicate higher accuracy.
From the heatmap, we observe that SR-GM has robust performance across most configurations. A notable exception occurs at $r$ of 0.05%, where performance degrades for $\lambda$ values of 500 and 1000. It shows that while SR-GM is insensitive to high values of $\lambda$, a correlation still exists between $r$ and $\lambda$. Specifically, we should use a larger $\lambda$ when $r$ increases to maintain optimal performance.
A.5. Proof of Theorem 2.1
Theorem 2.1.
Under gradient matching optimization, the amplification factor of modal mixing noise propagated through the graph structure is bounded by the Dirichlet energy of the gradient field:
$\|\mathbf{L}_{S}\mathbf{R}\|_{F}^{2}\leq\lambda_{\max}(\mathbf{L}_{S})\cdot E(\mathbf{R}),$
where $E(\mathbf{R})=\operatorname{tr}(\mathbf{R}^{\top}\mathbf{L}_{S}\mathbf{R})$ is the Dirichlet energy of the gradient field, and $\lambda_{\max}(\mathbf{L}_{S})$ is the maximum eigenvalue of the Laplacian matrix.
Proof.
We first define the gradient noise matrix $\mathbf{R}\in\mathbb{R}^{N^{\prime}\times d}$, where each row $\mathbf{r}_{i}^{\top}$ represents the gradient noise of node $i$. The graph Laplacian matrix $\mathbf{L}_{S}\in\mathbb{R}^{N^{\prime}\times N^{\prime}}$ is symmetric positive semi-definite and can be decomposed via eigenvalue decomposition:
$\mathbf{L}_{S}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top},$
where $\mathbf{U}$ is an orthogonal matrix, $\boldsymbol{\Lambda}=\operatorname{diag}(\lambda_{1},\lambda_{2},\dots,\lambda_{N^{\prime}})$, and the eigenvalues satisfy $0=\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{N^{\prime}}=\lambda_{\max}(\mathbf{L}_{S})$.
After propagation through the Laplacian matrix, the noise becomes $\mathbf{L}_{S}\mathbf{R}$, and the square of its Frobenius norm is:
$\|\mathbf{L}_{S}\mathbf{R}\|_{F}^{2}=\operatorname{tr}\left(\mathbf{R}^{\top}\mathbf{L}_{S}^{\top}\mathbf{L}_{S}\mathbf{R}\right)=\operatorname{tr}\left(\mathbf{R}^{\top}\mathbf{L}_{S}^{2}\mathbf{R}\right).$
Substituting the eigenvalue decomposition:
[…]
where $\tilde{\mathbf{R}}=\mathbf{U}^{\top}\mathbf{R}$. Let $\tilde{\mathbf{r}}_{i}^{\top}$ be the $i$-th row of $\tilde{\mathbf{R}}$, then:
$\|\mathbf{L}_{S}\mathbf{R}\|_{F}^{2}=\sum_{i=1}^{N^{\prime}}\lambda_{i}^{2}\|\tilde{\mathbf{r}}_{i}\|^{2}.$
On the other hand, the Dirichlet energy is defined as:
$E(\mathbf{R})=\operatorname{tr}\left(\mathbf{R}^{\top}\mathbf{L}_{S}\mathbf{R}\right)=\operatorname{tr}\left(\mathbf{R}^{\top}\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}\mathbf{R}\right)=\operatorname{tr}\left(\tilde{\mathbf{R}}^{\top}\boldsymbol{\Lambda}\tilde{\mathbf{R}}\right)=\sum_{i=1}^{N^{\prime}}\lambda_{i}\|\tilde{\mathbf{r}}_{i}\|^{2}.$
Since all eigenvalues are non-negative and satisfy $\lambda_{i}\leq\lambda_{\max}$, we have $\lambda_{i}^{2}\leq\lambda_{\max}\lambda_{i}$, thus:
$\sum_{i=1}^{N^{\prime}}\lambda_{i}^{2}\|\tilde{\mathbf{r}}_{i}\|^{2}\leq\lambda_{\max}\sum_{i=1}^{N^{\prime}}\lambda_{i}\|\tilde{\mathbf{r}}_{i}\|^{2},$
which is equivalent to:
$\|\mathbf{L}_{S}\mathbf{R}\|_{F}^{2}\leq\lambda_{\max}(\mathbf{L}_{S})\cdot E(\mathbf{R}).$
Equality holds if and only if all non-zero components of $\mathbf{R}$ lie in the eigenspace corresponding to the maximum eigenvalue. This inequality demonstrates that the amplification of gradient noise by the graph structure is bounded by the maximum eigenvalue of the Laplacian matrix and the Dirichlet energy of the gradient field.